• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Camsonne

    Log entry time 14:00:33 on June 12, 2010

    Entry number 326038

    keyword=Cavity monitor calibration attempt

    We tried with Ahmed to set the gains to get the Cavity monitor to work during the Q2 measurement. But it seems we could not see the beam at current less than 10 nA. It worked fine at 50 nA. According to Ahmed it did not work with higher gain and it seems we are very close from saturation at 0.5 uA. So maybe next time we should try the cross check of BPM and cavities at 0.3 uA and maybe go to the nA level in 2 steps. I will check